OpenStudy (anonymous):

The pilots of an airline have signed a new two-year contract. For the first year of the new contract, current salaries will be reduced by 25 percent. For the second year, these reduced salaries will be further reduced by 10 percent. What percent of their current salaries will the pilots earn during the second year of the new contract? Can anyone show me how to solve this, please?

OpenStudy (anonymous):

let the current salary be x it is reduced by 25 % so the salary for the first yr is 75 % of x now for the second year the salary is reduced by 10% from the salary of the first year . which means the salary of the second year is : 75% of x - 10% of 75% of x which is 90% of 75 % of x that is (90/100) *(75/100)*x = 67.5 % of x there the salary for second year is 67.5% of current salary

OpenStudy (anonymous):

current salary = y; \[y_1=0.75y\\ y_2=0.9y_1=0.9*0.75y=0.675y=67.5%y \]